Find lim h→0 f(h+3)-f(3)/h

a. f(x) = x²-3

b. f(x) = x/x-2

c. f(x) = sqrt x+1


1
Expert's answer
2021-11-03T18:02:54-0400

a. f(x)=x2-3

f(h+3)=(h+3)2-3=h2+6h+9-3=h2+6h+6

f(3)=32-3=9-3=6

"\lim\limits_{h\rarr 0}\frac{f(h+3)-f(3)}{h}=\lim\limits_{h\rarr 0}\frac{(h^2+6h+6)-(6)}{h}"

"=\lim\limits_{h\rarr 0}\frac{h^2+6h+6-6}{h}"

"=\lim\limits_{h\rarr 0}\frac{h^2+6h}{h}"

"=\lim\limits_{h\rarr 0}\frac{h(h+6)}{h}"

"=\lim\limits_{h\rarr 0}{h+6}"

=0+6

=6


b. f(x)="\frac{x}{x-2}"

f(h+3)="\frac{h+3}{h+3-2}=\frac{h+3}{h+1}"

f(3)="\frac{3}{3-2}=\frac{3}{1}=3"

"\lim\limits_{h\rarr 0}\frac{f(h+3)-f(3)}{h}=\lim\limits_{h\rarr 0}{\frac{\frac{h+3}{h+1}-3}{h}}"

"\lim\limits_{h\rarr 0}{\frac{\frac{h+3}{h+1}-3}{h}}=\lim\limits_{h\rarr 0}{\frac{\frac{h+3-3(h+1)}{h+1}}{h}}"

="\lim\limits_{h\rarr 0}{\frac{\frac{h+3-3h-3}{h+1}}{h}}"

="\lim\limits_{h\rarr 0}{\frac{\frac{h-3h}{h+1}}{h}}=\lim\limits_{h\rarr 0}({{\frac{h-3h}{h+1}}*\frac{1}{h}})"

="\lim\limits_{h\rarr 0}{{\frac{h-3h}{(h+1)h}}}=\lim\limits_{h\rarr 0}{{\frac{h(1-3)}{h(h+1)}}}"

"=\lim\limits_{h\rarr 0}{{\frac{1-3}{h+1}}}"

"=\frac{-2}{0+1}=\frac{-2}{1}"

=-2


c. f(x)="\sqrt{x+1}"

f(h+3)="\sqrt{h+3+1}=\sqrt{h+4}"

f(3)="\sqrt{3+1}=\sqrt{4}=2"

"\lim\limits_{h\rarr 0}\frac{f(h+3)-f(3)}{h}=\lim\limits_{h\rarr 0}\frac{\sqrt{h+4}-2}{h}"

"\lim\limits_{h\rarr 0}\frac{\sqrt{h+4}-2}{h}=\lim\limits_{h\rarr 0}(\frac{\sqrt{h+4}-2}{h}*\frac{\sqrt{h+4}+2}{\sqrt{h+4}+2})"

"=\lim\limits_{h\rarr 0}\frac{h+4+2\sqrt{h+4}-2\sqrt{h+4}-4}{h(\sqrt{h+4}+2)}"

"=\lim\limits_{h\rarr 0}\frac{h}{h(\sqrt{h+4}+2)}=\lim\limits_{h\rarr 0}\frac{1}{\sqrt{h+4}+2}"

"=\frac{1}{\sqrt{0+4}+2}=\frac{1}{\sqrt{4}+2}"

"=\frac{1}{2+2}"

="\frac{1}{4}"